From: ank@apo.nmsu.edu Subject: 2.5m obslog 20010131 (MJD 51941) Date: Thu, 1 Feb 2001 06:43:41 -0700 (MST) Wednesday Jan 31 2001 Apache Point Observatory SDSS 2.5m Observing Log (MJD 51941) OBSERVING TEAM: =============== Atsuko Nitta Kleinman (APO) Jurek Krzesinski (APO) Scot Kleinman (APO) Bing Chen (JHU) PT VISITORS: ========= THE PLAN: ========= Science SYNOPSIS FOR NIGHT: =================== Opened:19:30 1:15 Closed:22:15 6:00 Spectroscopy during the 2nd half of the night. We had to use "fk5InFiber" for all plates tonight before we could find the field. 2 plates were declared done. Atsuko survived the first year at APO -- Feb. 1 is the work@APO anniversary for her. TELESCOPE STATUS: ================= Opened:19:30 1:15 Closed:22:50 6:00 Cart 7 mounted. DIMM: ===== INTERLOCKS: =========== none bypassed SOFTWARE USED: ============== IOP/SOP: v3_25_4 Watcher: v2_8_14 MCP: v4_4_2 TPM: v2_2_1 AstroDa: 14_34 TCC: 2.4.1 WEATHER SYNOPSIS (see below for hourly data): ============================================= Partly cloudy and high humidity around sunset. Humidity dropped when the wind shifted -- but cirrus remained most of the second half of the night. GENERAL COMMENTS (see Problems below for details): ================================================== We opened around 19:30, there were scattered clouds and moon. So we started with cart 7. BTW, Cart1 and 2 has expired on Jan. 29 and needs to be replugged. We should have asked them to be replugged. We will do so tonight. The initial pointing was rather poor and we only saw the tail of the fk5 star with 5s exposure. Once we centered the fk5 and issued stickyOffset, we were able to acquire the field easily. The telescope changed its focus rather fast as we were setting up. The image of the guide stars we were seeing was quite horrible looking shortly after focusing the telescope. They looked elongated. We were worried that maybe the M2 Force Limit Fault message was actually true --- but after checking mirror numbers and doing another adjustment of focus, the stars no longer looked coma-like. We believe the focus was changing so fast that even after the first focusing, the telescope moved out of focus quickly enough that we saw out-of-focus images. While we were setting up, the cart 7 optimal time has passed. But since we had no un-expired plates until 22:30, we decided to go on with setting up -- this also helped as training for Jurek who drove tonight. The S/N from one exposure from this plate is so very horrible (1.3 to 0.5) that we did not put it in the database log. We went to the next plate, cart 3. But right after we put the cart on, the humidity rose to closing condition. We reopened after the humidity dropped. There were cirrus so we decided to continue spectro observing with cart 5. When using f4log command, we got imager as the instrument in use when it is actually spectro. "show inst" in TCC showed "spectro" as it should, but f4log showed "imager". A copy of one output of f4log is in the problem section. We finished two plates -- 272, 290 -- and declared them done. Both took 4 exposures to finish. The wind was remarkably calm all night long, but around morning twilight, the wind started become strong gain (20mph). The screw on the hand paddle (west side) came off. Perhaps the thread has worn off. We cannot hang the paddle without the screw, so we left the hand paddle on the garbage can. A PR was filed. endNIght ran successfully DATA ACQUISITION (see IOP/SOP output below): =========================================== IMAGING: -------- Run Type --- ---- 2091 bias in dog house --> one frame is saved in /data/scratch/ank SPECTROSCOPY: ------------- Exposure Type -------- ---- cart 7 plate 412 8472 bias 8473 preflat 8474 prearc 8475 science 8476 smear 8477 postflat 8478 postarc cart 5 plate 272 8479 preflat 8480 prearc 8481 science 8482 science 8483 science 8484 smear 8485 science 8486 postflat 8487 postarc cart 9 plate 290 8488 preflat 8489 prearc 8490 science 8491 science 8492 science 8493 science 8394 smear 8395 postflat 8396 postarc MIRRORS: ======== PRIMARY: -------- MIGS SHOULD BE Axial A=0.0370 0.015 Axial B=-0.0469 -0.051 Axial C=0.0299 0.030 Transverse D=3.7590 3.7569 Lateral E=1.90791 1.9104 Lateral F=0.4090 0.4095 GALILS Commanded: 4900. -2750. 450. -6650. 8750. 8300. Actual: -41600. -49250. -46000. -6650. 8950. 8500. SETMIR VALUES primDesOrient: 0.00 -4.40 10.70 586.50 240.70 primOrient: 0.00 -4.56 10.53 586.58 240.03 SECONDARY: ---------- Focus: 100.0 Air temp: -6.7 C Altitude: 56.77 MIGS A=-1.2949 B=-1.2949 C=-1.3070 D=0.2829 GALILS Commanded: 1589107. 1589107. 1589107. 7724. 8924. Actual: 4939809. 4937806. 4942134. 2150. 3350. SETMIR VALUES secDesOrient: 1260.87 0.00 0.00 0.00 -292.64 secOrient: 1260.93 0.40 -0.11 -0.07 -292.35 FOCUS LOG: ========== Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir fwhm 02:19 spectr 0.99990 988 2080 100 -18 56.0 -6.7 9 352 2.4 04:10 spectr 0.99990 988 1978 0 -9 56.9 -6.6 8 344 2.6 LN2 SCHEDULE: ============= We swapped imager dewar while we were closed due to high humidity. imager@206.8 spectr@150.8 with autofill connected PROBLEMS: ========= Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm 02:19 imager 0.99990 988 2080 100 -18 56.0 -6.7 9 352 ^ | This should have been spectro WEATHER DATA: ============= Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky Precip. 17:01 22 18 5 196 (SSW) 331 pcl 18:01 19 16 5 176 (S) 353 pcl 19:01 20 9 4 180 (S) 457 pcl 20:01 19 13 4 184 (S) 664 clr 21:01 18 11 8 190 (S) 715 clr 22:01 19 15 7 216 (SW) 800 clr 23:01 17 15 7 218 (SW) 794 clr 00:01 17 13 3 264 (W) 752 clr 01:01 20 7 1 263 (W) 713 cirrus 02:01 20 1 8 349 (NNW) 638 clr 03:01 20 0 8 354 (N) 199 cirrus 04:01 19 2 8 346 (NNW) 211 cirrus 05:01 20 1 4 48 (NE) 197 clr [pmt]cl = partly, mostly, totally cloudy DATA LOGS FROM IOP/SOP: =======================